OpenStudy (anonymous):

what is tan theta squared minus one factored? than (squared) theta - 1 (I need this factored)

OpenStudy (anonymous):

*tan

OpenStudy (solomonzelman):

use \[USE~~THIS:~~~a^2-b^2=(a+b)(a-b)\]\[(Tanx)^2-1^2=?\]

OpenStudy (isaiah.feynman):